Circular No. 8375

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dg A. Vagnozzi, Stroncone, Italy, reports the discovery by D. De Pasquale,

F. Guerri, G. Guerri, M. Cristofanelli, S. Romanelli, and himself of an apparent supernova on unfiltered CCD images taken with 0.50-m and 0.30m reflectors on July 19.854 (at mag 17.1) and 20.841 UT (mag 17.0). The new object is located at α = 14^{h}59^{m}58^{s}.96, δ = +1^{o}53′25′′.6 (equinox 2000.0; average of measurements by S. Valentini from two images), which is ≈ 19′′ west and 3′′ south of the center of NGC 5806. H. Yamaoka, Kyushu University, reports that K. Itagaki (Teppo-cho, Yamagata, Japan) found SN 2004dg at mag 16.8 on an unfiltered CCD image taken on July 21.534, adding that nothing is visible at this location on his archival images of this galaxy (the most recent one taken in early June).

SUPERNOVA 2004dh IN MCG +04-1-48 Further to IAUC 8373, M. Moore and W. Li report the LOSS discovery

of an apparent supernova on an unfiltered KAIT image taken on July 21.47 UT (mag ∼ 17.5). The new object was confirmed on an earlier image taken on July 11.45 (mag ∼ 18.1). SN 2004dh is located at α = 0^{h}18^{m}00^{s}.39, δ = +24^{o}33′44′′.5 (equinox 2000.0), which is 7′′.3 east and 2′′.1 south of the nucleus of MCG +04-1-48. A KAIT image taken on June 30.48 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2004dd IN NGC 124 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 340–730 nm) of SN 2004dd (cf. IAUC 8373, 8374), obtained by M. Calkins on July 18.44 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-II supernova. Adopting the NED recession velocity of 4060 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 6300 km/s.

COMET C/2001 Q4 (NEAT) Visual total-magnitude estimates: June 13.91 UT, 6.4 (A. Diepvens,

Balen, Belgium, 20×50 binoculars); 30.11, 7.0 (J. J. Gonzalez, Leon, Spain, 7×50 binoculars); July 5.94, 7.3 (K. Hornoch, Lelekovice, Czech Republic, 10×50 binoculars); 12.91, 7.6 (A. Baransky, Bucha, Ukraine, 10×50 binoculars); 17.88, 7.8 (R. Haver, Frasso Sabino, Italy, 15×80 binoculars); 20.48, 7.9 (Y. Nagai, Nagano, Japan, 11×80 binoculars).

2004 July 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green